# Towards a Method for LLM-enabled Indoor Navigation (Arxiv) thanks: This work was founded by European Union - Next Generation EU, in the context of The National Recovery and Resilience Plan, Investment Partenariato Esteso PE8 ”Conseguenze e sfide dell’invecchiamento”, Project Age-IT, CUP: B83C22004880006

### v2

@@ Abstract
Indoor navigation presents unique challenges due to complex layouts, lack of GPS signals, and accessibility concerns. Existing solutions often struggle with real-time adaptability and user-specific needs. In this work, we explore the potential of a Large Language Model (LLM), i.e., ChatGPT, to generate natural, context-aware navigation instructions from indoor map images. We design and evaluate test cases across different real-world environments, analyzing the effectiveness of LLMs in interpreting spatial layouts, handling user constraints, and planning efficient routes. Our findings demonstrate the potential of LLMs for supporting personalized indoor navigation, with an average of 50.54% correct indications and a maximum of 77.78%. The results do not appear to depend on the complexity of the layout or the complexity of the expected path, but rather on the number of points of interest and the abundance of visual information, which negatively affect the performance.
@@
Inclusivity, Indoor localization, Large Language Model, AI-driven path planning, Accessibility, spatial reasoning, indoor maps.
@@ I Introduction
Efficient indoor navigation and scene representation are essential for various applications, including robotics, smart environments, and assistive technologies. Unlike outdoor navigation, which benefits from well-established global positioning systems (GPS) and extensive mapping databases, indoor navigation presents unique challenges [1]. The complexity and variability of enclosed spaces, the absence of GPS signals, and the need for detailed real-time scene understanding make indoor navigation a demanding task [2]. Current solutions rely on advanced computational models to interpret the environment and adapt to dynamic conditions [3]. However, they often lack flexibility, accessibility, and adaptability for diverse user needs.
There is a growing need for a system capable of providing personalized, natural-language navigation assistance to address these limitations[4]. Such a system should interpret indoor maps and generate human-like, context-aware navigation instructions tailored to individual users. A system like this would also be highly beneficial for visually impaired individuals, as a detailed, natural-language description of the environment can enhance spatial awareness, helping them build a mental representation of their surroundings and navigate with greater confidence and independence. Integrating multimodal inputs, such as textual queries, voice commands, or sensor data, could provide more adaptive and accessible navigation experiences, ultimately fostering greater independence for individuals facing navigational barriers[5]. Beyond accessibility challenges, navigating unfamiliar indoor environments is another scenario where such a system would be beneficial. People visiting a new location, such as airports, shopping malls, hospitals, or office buildings, often struggle with orientation, requiring clear and accessible guidance to reach their destinations efficiently. A common approach to overcoming these challenges is to ask other people for directions, who typically provide landmark-based instructions (e.g., you will see a shoe shop, turn right, then you will encounter a phone charging area…). Given these challenges, leveraging Large Language Models (LLMs) for indoor navigation presents a promising alternative. LLMs can interpret complex multimodal inputs, reason over spatial relationships, and generate human-like natural language indications tailored to user-specific queries. This adaptability makes them well-suited for generating dynamic, human-like navigation instructions based on indoor maps[6]. In fact, recent advances in LLMs have opened new avenues for user-friendly navigation solutions for indoor and outdoor contexts, and preliminary studies have been performed leveraging LLMs to address this goal. For example, in [7], authors introduce TraveLLM, a system for public transit route planning. Their study compares multiple LLMs—including GPT-4, Claude 3, and Gemini—demonstrating that these models can integrate real-time data and user-specific constraints to form contextually aware navigation recommendations than conventional platforms (e.g., Google Maps). However, their study focuses on outdoor scenarios, leaving indoor navigation unexplored. Meanwhile, Zhang et al. [8] focus on improving navigation for people with visual impairments (PVI). NaviGPT integrates LiDAR-based obstacle detection, vibration feedback, and LLM-generated instructions to produce a real-time user experience. Unlike existing tools, e.g., Be My AI [^1] https://www.seeingai.com/ and Seeing AI  [^2] https://www.bemyeyes.com/blog/introducing-be-my-ai, which require multiple apps for obstacle detection, image recognition, or contextual directions, NaviGPT consolidates these capabilities into one cohesive system. This seamless integration offers continuous feedback without user distraction from frequent application switching. However, no evaluation of the system has been provided, and at the current stage, the proposal appears to mainly target outdoor environments. For indoor context, in [9], authors examine the development of geo-descriptions—textual narratives that describe the internal layout of a building. These authors combine classical GIS spatial analysis and LLM-based text generation to automate the creation of human-readable indoor maps, which is traditionally labour-intensive. However, the authors do not consider querying the system for navigation paths in the reconstructed textual representation of the building. Furthermore, the evaluation is performed on simplified cases.
@@
Despite these preliminary studies, the use of visual maps as input for LLMs to provide indications based on user requests has not been explored. This could highly simplify the architecture of indoor navigation systems, thus making them easily adaptable to different indoor environments. Once a visual map is made available—as common for airports, shopping malls, and large environments with several points of interest—a system based on these principles is expected to be ready to use in any environment without complex, ad hoc, configurations. Indeed, this approach avoids the task of providing structured representations of the maps for each environment, which can be time-consuming.
To address the research gap, we evaluate the potential of an LLM, namely GPT-4, to assist in indoor navigation using three real-world interior 2D maps of buildings, including two airport terminal maps and one shopping mall map. We investigate the LLM ability to generate step-by-step guidance based on diverse user preferences and spatial constraints. A usage scenario is depicted in Fig. 1, where a user asks the LLM for directions to a specific gate and receives human-like indications to navigate the environment. Using different real scenarios, we want to evaluate different test cases across three real-world environments, analyzing the effectiveness of LLMs in interpreting spatial layouts, handling user constraints, and planning efficient routes.
@@
    We propose an LLM-based approach for generating personalized navigation instructions based on indoor maps.
@@
    We evaluate the effectiveness of GPT-4 in generating accurate, human-like navigation guidance.
By exploring the role of LLMs in indoor navigation, this work aims to bridge the gap between traditional map-based systems and intelligent, real-time, and user-adaptive guidance solutions.
Figure 1: Example of an LLM assisting with indoor navigation tasks.
All supplementary material, including prompts, input maps, and results, are publicly available  [^3] https://doi.org/10.5281/zenodo.14993277.
II Understanding user needs: Requirements for an inclusive experience
Today, many public venues have navigational aids like digital signage, online platforms, or smartphone apps. However, these aids frequently fail to accommodate individuals with different requirements, and the general public often encounters difficulties due to unclear or incorrect directions.
For instance, mobile apps that use digital maps often do not include features to assist people with visual impairments [10] or mobility disabilities. Nonetheless, navigation apps designed for accessibility have shown to be beneficial for such users [11]. A detailed and accurate description of the surrounding environment can enhance spatial awareness and improve the overall navigation experience. This is especially useful in complex indoor spaces, where clear, context-aware information can help users confidently reach their destinations. Designing navigation systems with adaptability in mind ensures they can satisfy a diverse range of needs, from visitors unfamiliar with a building’s layout to individuals who benefit from alternative navigation methods.
Indoor spaces vary widely in size, layout, and functionality, encompassing different types of destinations such as stores, restrooms, and emergency exits. Users often ask questions such as “What can I find on this floor?” or “Where is the nearest restroom?” highlighting the need for adaptable navigation solutions that can dynamically respond to different environments and user preferences. Among the many scenarios where such systems can be particularly beneficial is assisting users who may face additional navigation challenges, such as visual or hearing impairments, mobility constraints, or cognitive differences. While many existing systems are not designed to accommodate a broad spectrum of interaction needs, a more advanced approach could integrate personalized, multimodal assistance, ensuring that users receive navigation guidance in a format that best suits their preferences.
@@ II Understanding user needs: Requirements for an inclusive experience
Particular attention must be paid to the user experience. Navigating a complex indoor environment can be challenging for anyone, whether due to unfamiliarity with the building’s layout, unclear signage, or language barriers. These difficulties often lead users to seek assistance, emphasizing the need for a system that provides clear, accessible, and efficient guidance. To be truly effective, such a system must offer a seamless and intuitive experience. In our vision of a functional and efficient solution, users should be able to input personal preferences or specific needs and receive tailored navigation assistance in real-time. The most effective way to achieve this is through natural language interaction.
Natural language is not only a valid and immediate input method but also makes the system accessible to many users. The system can accommodate diverse needs by allowing users to interact through written or spoken text, ensuring a more inclusive and user-friendly navigation experience.
@@ III Experimental Set-up
This section illustrates the approach we followed to evaluate the capability of the LLM to support indoor navigation. Given a user query, we expect the LLM to deliver a detailed output, including step-by-step directions, points of interest, and potential obstacles, all aimed at enhancing the user’s navigation experience.
Figure 2 describes the pipeline of our evaluation. Its four main components are:
  • Image Preprocessing Initially, we manually removed the legend from the image, to ensure that the input map is free of noise that can affect the LLM performance. This step was introduced after preliminary tests, in which we identified several errors likely caused by this information. It should be noted that this step is performed once for each used map, and not at runtime.
  • System prompt This prompt is hard-coded for the LLM and will be used for any user query. It includes all the necessary instructions, details, and potential errors to avoid. This prompt is the result of several iterations through which it has been improved to prevent critical and systematic errors.
  • User Output The system produces an output with human-like indications. This output can be delivered in written or spoken form, depending on the user preferences and possible accessibility requirements.
Certain potential errors were identified and avoided during the prompt engineering phase. These errors included the model misinterpreting points that are close together on the map legend as being geographically proximate and points that appear numerically adjacent. Moreover, an overabundance of information was initially provided as output, which could overwhelm the user. A uniform approach was adopted to address these issues where a single user prompt structure—with different content—was used across all tests. This format does not vary between maps or include detailed descriptions specific to any map. Instead, references to locations in the user prompts are made either by the numerical ID of a point of interest or by the point’s name, depending on the map’s representation of the points. It should be highlighted that, in a practical scenario with a real user querying the LLM, prompt engineering is not feasible. In future work, we will study how to translate natural user queries into engineered prompts to better exploit the LLM’s potential.
The tests were conducted using three different maps: the ground floor plans of Bergamo Airport, Bologna Airport, and Orio Center—a shopping mall. These maps were sourced from the facilities’ official websites. The images differ from each other and aim to encompass various structural scenarios. The airports are characterized by a main corridor where various points of interest for guests are located on the sides, featuring many services and different types of access. Bologna Airport has a less dense set of points of interest than Bergamo Airport. In contrast, the Orio Center is characterized by a denser network of corridors, with various shops and points of interest.
@@
For each map, five requests are made, each with a different category of difficulty. These were defined ad hoc based on a visual analysis of each map. The first three queries differ in terms of distance between starting and destination points, i.e., short (S), medium (M), and long (L). The other two queries, referred to as “specific requests” (Spec), explore other elements of complexity, which depend on each map. Specifically, we challenged the LLM with routes that have hidden or poorly marked IDs of points of interest, or with prompts in which the specific destination was not explicitly stated, e.g., “Show me how to walk to the nearest dining area from gate number 26”. In this case, the query requires a more detailed search near gate 26 to locate the closest dining area. Table III reports examples of user prompts for the Bologna Airport.
@@
The evaluation of the LLM’s responses is carried out by defining a ground truth of correct paths. Once the start and destination points are established for each query, we manually identify the optimal route by imagining a user standing at the starting point and then considering the shortest, least convoluted path through the structure. If there are multiple routes of similar length, they are all considered correct.
@@
  • Incorrect – the instruction points to a different direction with respect to the destination, or asks to follow routes that are not physically possible, e.g., across walls;
Figure 3: Correct path considered for ORIO-2 query (Table  IV).
Let us consider one example to exemplify the correct and incorrect cases and the rationale for our tagging. Taking the query “How to go from Piazza Italia to Milani Cafè”—cf. Figure 3 for the area of interest and the correct path—this yields the following output:

### v4

@@ Abstract
Indoor navigation presents unique challenges due to complex layouts and the unavailability of GNSS signals. Existing solutions often struggle with contextual adaptation, and typically require dedicated hardware. In this work, we explore the potential of a Large Language Model (LLM), i.e., ChatGPT, to generate natural, context-aware navigation instructions from indoor map images. We design and evaluate test cases across different real-world environments, analyzing the effectiveness of LLMs in interpreting spatial layouts, handling user constraints, and planning efficient routes. Our findings demonstrate the potential of LLMs for supporting personalized indoor navigation, with an average of 86.59% correct indications and a maximum of 97.14%. The proposed system achieves high accuracy and reasoning performance. These results have key implications for AI-driven navigation and assistive technologies.
@@
Large Language Model, Multimodal, AI-driven path planning, Indoor Navigation, Accessibility, Spatial Reasoning, Indoor Maps.
@@ I Introduction
Efficient indoor navigation and scene representation are essential for various applications, including robotics, smart environments, and assistive technologies. Unlike outdoor navigation, which benefits from well-established global positioning systems (GPS) and extensive mapping databases, indoor navigation presents unique challenges [1]. The complexity and variability of enclosed spaces, the absence of GPS signals, and the need for detailed real-time scene understanding make indoor navigation a demanding task [2].
Current solutions rely on advanced computational models to interpret the environment and adapt to dynamic conditions [3]. However, they often lack flexibility, generalization capability, and require the support of dedicated hardware.
There is a growing need for a system capable of providing personalized natural language navigation assistance to address these limitations[4]. Such a system should interpret indoor maps and generate human-like, context-aware navigation instructions tailored to individual users.
People visiting a new location, such as airports, shopping malls, hospitals, or office buildings, often struggle with orientation, requiring clear and accessible guidance to reach their destinations efficiently. A common approach to overcoming these challenges is to ask other people for directions, who typically provide landmark-based instructions (e.g., you will see a shoe shop, turn right, then you will encounter a phone charging area…).
Given these challenges, leveraging Large Language Models (LLMs) for indoor navigation presents a promising alternative. LLMs can interpret complex multimodal inputs, reason over spatial relationships, and generate human-like natural language indications tailored to user-specific queries. This adaptability makes them well-suited for generating dynamic, human-like navigation instructions based on indoor maps[5].
In fact, recent advances in LLMs have opened new avenues for user-friendly navigation solutions for indoor and outdoor contexts, and preliminary studies have been performed leveraging LLMs to address this goal.
For example, in [6], the authors introduce TraveLLM, a system for public transit route planning. Their study compares multiple LLMs—including ChatGPT-4, Claude 3, and Gemini—demonstrating that these models can integrate real-time data and user-specific constraints to form contextually aware navigation recommendations than conventional platforms (e.g., Google Maps). However, their study focuses on outdoor scenarios, leaving indoor navigation unexplored.
Meanwhile, Zhang et al. [7] focus on improving navigation for people with visual impairments (PVI). NaviGPT integrates LiDAR-based obstacle detection, vibration feedback, and LLM-generated instructions to produce a real-time user experience. Unlike existing tools, e.g., Be My AI [^1] https://www.bemyeyes.com/blog/introducing-be-my-ai and Seeing AI [^2] https://www.seeingai.com/, which require multiple apps for obstacle detection, image recognition, or contextual directions, NaviGPT consolidates these capabilities into one cohesive system. This seamless integration offers continuous feedback without user distraction from frequent application switching. However, no evaluation of the system has been provided, and at the current stage, the proposal appears to mainly target outdoor environments.
For indoor context, in [8], the authors examine the development of geo-descriptions—textual narratives that describe the internal layout of a building. These authors combine classical GIS spatial analysis and LLM-based text generation to automate the creation of human-readable indoor maps, which is traditionally labour-intensive. However, the authors do not consider querying the system for navigation paths in the reconstructed textual representation of the building. Furthermore, the evaluation is performed on simplified cases.
@@
Despite these preliminary studies, the use of visual maps as input for LLMs to provide indications based on user requests has not been explored. This could highly simplify the architecture of indoor navigation systems, thus making them easily adaptable to different indoor environments. Once a visual map is made available—as is common for airports, shopping malls, and large environments with several points of interest—a system based on these principles is expected to be ready to use in any environment without complex, ad hoc, configurations. Indeed, this approach avoids the task of providing structured representations of the maps for each environment, which can be time-consuming.
To address the research gap, we evaluate the potential of an LLM, namely ChatGPT-4 (in two versions, ChatGPT-4o and the superior o3), to assist in indoor navigation using three real-world interior 2D maps of buildings, including two airport terminal maps and one shopping mall map.
We investigate the LLM ability to generate step-by-step guidance based on diverse user preferences and spatial constraints.
Using different real scenarios, we want to evaluate different test cases across three real-world environments, analyzing the effectiveness of LLMs in interpreting spatial layouts, handling user constraints, and planning efficient routes.
@@
    We evaluate the effectiveness of ChatGPT-4 in generating accurate, human-like navigation guidance based on visual indoor maps.
@@
    According to [9], all supplementary materials, including prompts, input maps, annotation guide, and results, are publicly available [^3] https://doi.org/10.5281/zenodo.14993277.
This work explores the role of LLMs in indoor navigation, aiming to bridge the gap between traditional map-based systems and intelligent, real-time, and user-adaptive guidance solutions.
@@
II System Overview and Pipeline Description
This section illustrates the approach we followed to evaluate the capability of the LLM to support indoor navigation. Given a user query, we expect the LLM to deliver a detailed output, including step-by-step directions, points of interest (POIs), and potential obstacles, all aimed at enhancing the user’s navigation experience.
We performed three different experiments, considering five different types of queries, and three maps. Each experiment corresponds to an incremental refinement, e.g., different pre-processing steps or model versions, introduced to overcome the limitations of the previous experiment. All the experiments are based on the same pipeline as described in the following.
Preliminary tests were conducted using several state-of-the-art LLM frontier models, including Gemini, ChatGPT, and Claude. Based on these initial evaluations, the model that delivered the most promising results—ChatGPT—was selected for the experiments presented in this study.
Figure 1 describes the pipeline of our experimental process. Its four main components are:
  • Map Pre-processing The input images were manually adjusted to enhance performance. A range of Pre-processing techniques (e.g., removal of the legend, simplification of the map) was applied throughout the different experiments, each informed by specific hypotheses or objectives. Detailed descriptions of the manual pre-processing steps are provided in each experiment analysis section, i.e., Section IV-A, IV-B, and IV-C.
  • System Prompt The system prompt is hard-coded for the language model and is applied uniformly to all user queries. It contains comprehensive instructions, relevant details, and guidelines for avoiding potential errors. This prompt is the result of several iterative refinements aimed at preventing critical and systematic mistakes, and enhancing the model’s overall performance. Across the various experiments, two different versions of the system prompt were employed. The specific version used in each experiment is disclosed and discussed in the corresponding experiment analysis sections.
  • User Output The system produces an output with human-like indications delivered in written form.
A single user prompt structure—with different content—was used across all images in an experiment. The structure does not vary between maps and does not include detailed descriptions specific to any map. Instead, references to locations in the user prompts are made either by the numerical ID of a point of interest or by the point’s name, depending on the map’s representation of the points. It should be highlighted that, in a practical scenario with a real user querying the LLM, prompt engineering is not feasible.
@@ III Experimental Set-up
The experiments were conducted using three different maps, all in high quality .png format:
  • Bergamo The Terminal map is visually clear and well-organized, with labeled POIs and a straightforward layout, but the high density and small size of elements hinder accurate image parsing by the language model.
  • Bologna Similar to Bergamo airport, this map has a central corridor flanked by shops, but it differs in its color scheme, floor representation, and a more curved, less linear layout.
  • Orio Center This shopping mall map was selected for its complex layout, featuring frequent directional changes and intersecting corridors, with small retail outlets embedded in the paths, making spatial interpretation more challenging.
@@
For each map, five requests are made, each with a different category of difficulty. These were defined ad hoc based on a visual analysis of each map. The first three queries differ in terms of distance between starting and destination points, i.e., short (S), medium (M), and long (L). The other two queries, referred to as “specific requests” (Spec), explore other elements of complexity, which depend on each map. Specifically, we challenged the LLM with routes that have hidden or poorly marked IDs of POIs, or with prompts in which the specific destination was not explicitly stated, e.g., “Show me how to walk to the nearest dining area from gate number 26”. In this case, the query requires a more detailed search near gate 26 to locate the closest dining area. Table I reports examples of user prompts for the Bologna Airport.
Each reference to query number X of a specific map will be cited as follows: BER-X for Bergamo airport, BOL-X for Bologna airport, and ORIO-X for Orio Center.
TABLE I: User prompt examples for the Bologna Airport.
Scenario ID Query Category User Prompt Query Rationale
III-A Evaluation Method
@@
The evaluation of the LLM’s responses is carried out by defining a ground truth of correct paths. Once the start and destination points are established for each query, we manually identify the optimal route by imagining a user standing at the starting point and then considering the shortest, least convoluted path through the structure.
@@
  • Incorrect – the instruction does not reflect the actual state of the navigation or provides misleading information to the user, such as indicating the wrong direction or suggesting physically impossible paths (e.g., through walls).
Figure 2: Correct path considered for ORIO-2 query (Table I) on the Orio Center mall map.
In the supplementary material, we share the annotation guide adopted to identify path segments and label them, to ensure the replicability of our experiments. In the following, we consider one example to showcase correct and incorrect cases and the rationale for our tagging. Taking the query “How to go from Piazza Italia to Milani Cafè”—cf. Figure 2 for the area of interest and the correct path—this yields the following output: